Exercise 11-5 (Video) The standard cost of Product B manufactured by Pharrell Company includes 2.0 units of direct materials at $6.90 per unit. During June, 27,500 units of direct materials are purchased at a cost of $6.85 per unit, and 27,500 units of direct materials are used to produce 13,700 units of Product B. (a) Compute the total materials variance and the price and quantity variances. Total materials variance $ Materials price variance $ Materials quantity variance $ (b) Compute the total materials variance and the price and quantity variances, assuming the purchase price is $6.95 and the quantity purchased and used is 26,500 units. Total materials variance $ Materials price variance $ Materials quantity variance $ Click if you would like to Show Work for this question: Open Show Work
Answer #1

Solution :

SP = 6.90
SQ = 27400 (13700*2)
AP = 6.85
AQ = 27500
Price Veriance is AQ(SP-AP) 27500 (6.90-6.85) = 1375 (This is favorable, the material
were obtained at below average cost)
Quantity Veriance SP (SQ-AQ) 6.90 (27400-27500) = 6.90 (-100) = -690 This is unfavorable
more than the standered quantity was used)
Total material variance can be obtained : SQ*SP- AQ*AP
27400 x 6.90 -27500 x 6.85
=685 (Favorable)
2
SP = 6.90
SQ = 27400 (13700 x 2)
AP = 6.95
AQ = 26500
Price Veriance : AQ (SP-AP) 26500 (6.90-6.95) = -1325 (This is unfavorable Veriance)
Quantity Veriance SP (SQ-AQ) 6.90 (27400-26500) = 6210 (This is favorable Veriance)
Total material variance can be obtained : SQ*SP- AQ*AP
27400 x 6.90- 26500 x 6.95
= 4885 (Favorable)